# Background independent tensor networks

### v2

@@ Abstract
Conventional holographic tensor networks can be described as toy holographic maps constructed from many small linear maps acting in a spatially local way, all connected together with “background entanglement”, i.e. links of a fixed state, often the maximally entangled state. However, these constructions fall short of modeling real holographic maps. One reason is that their “areas” are trivial, taking the same value for all states, unlike in gravity where the geometry is dynamical. Recently, new constructions have ameliorated this issue by adding degrees of freedom that “live on the links”. This makes areas non-trivial, equal to the background entanglement piece plus a new positive piece that depends on the state of the link degrees of freedom. Nevertheless, this still has the downside that there is background entanglement, and hence it only models relatively limited code subspaces in which every area has a definite minimum value. In this note, we simply point out that a version of these constructions goes one step further: they can be background independent, with no background entanglement in the holographic map. This is advantageous because it allows tensor networks to model holographic maps for larger code subspaces. In addition to pointing this out, we address some subtleties involved in making it work and point out a nice connection it offers to recent discussions of random CFT data.
@@ 1 Introduction
Let us summarize the network now. First, how might we make a tensor network without a rigid, fixed background? As a first guess, we might imagine a model in which we consider more than just one tensor network, allowing (somehow) for a “superposition of tensor networks”, each with a different geometry. This is a decent first step towards modeling gravity. However, it fails to model the fact that in gravity we do not allow arbitrary quantum states on arbitrary geometries. There are constraints that the geometry must satisfy. These constraints are important and related to having good time evolution matching that of the dual theory. Really, we would like to construct a model akin to a “superposition of tensor networks” but in which the geometries of the tensor networks are required to satisfy some constraints.
@@
That said, these existing constructions arguably only go part of the way. They involve a background geometry, in a manner we’ll explain. The main goal of this note is to point out that a version of these constructions goes all the way, continuing to work even without the background geometry, once we have incorporated link degrees of freedom in the appropriate way. We explain this in Section 4. Thus we arrive at our goal: a tensor network incorporating multiple geomtries, possibly all very distinct, and satisfying certain constraints. In Section 4.2 we address a subtlety in how to get this to work in one spatial dimension, and in Section 4.3 we point out that these background independent tensor networks allow for a straightforward connection to recent discussions on gravity-from-random-CFT-data.
@@ 3 Link degrees of freedom
Now we turn to the tensor network. We want something like the one in Section 2, but modified to act on this new $\mathcal{H}_{\mathrm{bulk}}$. The trick is we’ll first define a new kind of tensor that takes as input a link state in $\mathcal{H}_{G}$ and outputs a state in a factorized Hilbert space. Specifically, we define a tensor $S_{(xy)}:\mathcal{H}_{G}^{(xy)}\to\mathcal{H}_{G}^{xy}\otimes\mathcal{H}_{G}^{yx}$ to implement the following factorization (the usual embedding into the “extended Hilbert space” [23])
@@
with $\braket{\psi}{S^{\dagger}\hat{A}(b)S}{\psi}=\sum_{xy\in\partial b}\log D_{xy}$ (here for simplicity specializing to the case |$\ket{\phi_{xy}}=\ket{\mathrm{MAX}}$). This is the correct answer (again, in the limit all $D_{xy}$ are very large), but we would like the answer for $V$ instead of $\widetilde{V}$. Fortunately, by construction $S(B)_{V\ket{\psi}}=S(B)_{\widetilde{V}S\ket{\psi}}$! All that’s different is the interpretation of the terms on the right hand side of (3.16). For $V$, the factors $\mathcal{H}_{G}^{xy}$ are not part of $\mathcal{H}_{\mathrm{bulk}}$, and so their entropy – the third term in the last line of (3.15) – is not a part of $S(b)_{\ket{\psi}}$. Instead, those terms must be interpretted as part of the area term. We end up with
@@ 4.2 $1d$ and background independence
Now let’s try to compute the entropy $S(B)$ for $B$ either of the two black circles. All bond dimensions are equal, so the minimal $b$ in the formula (4.10) can have many problems. For example, it can be highly degenerate, say if the qudits are in a product state. More problematically, if any bulk legs are in a mixed state, then neither boundary region will include that leg in its minimal $b$. (Though the entire boundary might still include the entire bulk in its minimal $b$.) This is not like what we expect in $1$-dimensional versions of AdS/CFT.
@@
4.3 Random tensors as random Clebsch-Gordan coefficients
Having removed the background entanglement, we can now draw an interesting connection to recent ideas relating gravity to an ensemble of CFT data [26, 27]. [^13] We thank Ronak Soni for discussions on this. Consider a network from Section 4.1, with all trivalent vertices and no matter degrees of freedom. $\mathcal{H}_{\mathrm{bulk}}$ is just a lattice gauge theory. We’ll now show that the choice of random tensor admits an interpretation as a choice of “random Clebsch-Gordan coefficient”. This is analogous to random CFT data because in AdS/CFT, the bulk fusion coefficients relate to OPE coefficients in the dual CFT.
To see the random Clebsch-Gordan coefficients here, consider again the tensors $\bra{t_{x}}$ | from (4.5), this time with just the gauge theory and for a trivalent vertex. The (relevant part of the) tensor takes the form
$\bra{T^{\prime}_{x}}=\sum_{\mu,\mu^{\prime},\mu^{\prime\prime}\in M}\bra{t_{x}^{\mu,\mu^{\prime},\mu^{\prime\prime}}}\bra{\Omega_{\mu,\mu^{\prime},\mu^{\prime\prime}}}~{},$ (4.16)
where $M$ is the set of allowed $\mu,\mu^{\prime},\mu^{\prime\prime}$ fusions, and
$\bra{t_{x}^{\mu,\mu^{\prime},\mu^{\prime\prime}}}\in\mathcal{H}_{\mu}^{*}\otimes\mathcal{H}_{\mu^{\prime}}^{*}\otimes\mathcal{H}^{*}_{\mu^{\prime\prime}}$ (4.17)
is a random state that we can write as
$\bra{t_{x}^{\mu,\mu^{\prime},\mu^{\prime\prime}}}=\sum_{k,k^{\prime},k^{\prime\prime}}F^{\mu\mu^{\prime}\mu^{\prime\prime}}_{kk^{\prime}k^{\prime\prime}}\bra{\mu;k}\bra{\mu^{\prime};k^{\prime}}\bra{\mu^{\prime\prime};k^{\prime\prime}}$ | (4.18)
for random $F^{\mu\mu^{\prime}\mu^{\prime\prime}}_{kk^{\prime}k^{\prime\prime}}$, which we will ultimately interpret as the random Clebsch-Gordan coefficient. Recall $\bra{\Omega_{\mu,\mu^{\prime},\mu^{\prime\prime}}}\in\mathcal{H}_{\mu^{*}}^{*}\otimes\mathcal{H}_{{\mu^{\prime}}^{*}}^{*}\otimes\mathcal{H}^{*}_{{\mu^{\prime\prime}}^{*}}$ is a fixed state determined by the (true, non-random) Clebsch-Gordan coefficients $C^{\mu\mu^{\prime}\mu^{\prime\prime}}_{ii^{\prime}i^{\prime\prime}}$ of the gauge group $G$,
$\bra{\Omega_{\mu,\mu^{\prime},\mu^{\prime\prime}}}:=C^{\mu\mu^{\prime}\mu^{\prime\prime}}_{ii^{\prime}i^{\prime\prime}}\bra{\mu;i}\bra{\mu^{\prime};i^{\prime}}\bra{\mu^{\prime\prime};i^{\prime\prime}}~{}.$ | . (4.19)
Note that $\braket{\Omega_{\mu_{1},\mu^{\prime}_{1},\mu_{1}^{\prime\prime}}}{\Omega_{\mu_{2},\mu^{\prime}_{2},\mu^{\prime\prime}_{2}}}=\delta_{\mu_{1},\mu_{2}}\delta_{\mu^{\prime}_{1},\mu^{\prime}_{2}}\delta_{\mu^{\prime\prime}_{1},\mu^{\prime\prime}_{2}}$. This $\bra{T^{\prime}_{x}}$ | acts on the Hilbert space $\mathcal{H}_{G}\otimes\mathcal{H}_{G}\otimes\mathcal{H}_{G}$ of the three links adjacent to $x$ after the action of $S$. Because of Gauss’ law at the vertex, a basis for gauge invariant states of these links takes the form
[…]
Now we see that when we act with $\bra{T^{\prime}_{x}}$ |, we might as well be acting
$\bra{\widetilde{T}^{\prime}_{x}}:=\sum_{\mu,\mu^{\prime},\mu^{\prime\prime}\in M}\bra{t_{x}^{\mu,\mu^{\prime},\mu^{\prime\prime}}}$ | (4.21)
on
|$\ket{\widetilde{\psi}}:=\sum_{\mu,\mu^{\prime},\mu^{\prime\prime}\in M}\ket{\mu;k}\ket{\mu^{\prime};k^{\prime}}\ket{\mu^{\prime\prime};k^{\prime\prime}}~{}.$ (4.22)
That is, the only randomness here was the choice of randomness of $F^{\mu\mu^{\prime}\mu^{\prime\prime}}_{kk^{\prime}k^{\prime\prime}}$ in the $k,k^{\prime},k^{\prime\prime}$ indices, as a function of the allowed $\mu,\mu^{\prime},\mu^{\prime\prime}$ configurations. In other words, the choice of random tensor is equivalent to a choice of random Clebsch-Gordan coefficient.
Note that this is not true when there is background entanglement. In that case the tensors $\bra{T^{\prime}_{x}}$ | take the form
$\bra{T^{\prime}_{x}}=\sum_{\mu,\mu^{\prime},\mu^{\prime\prime}\in M}\bra{p_{x}^{\mu,\mu^{\prime},\mu^{\prime\prime}}}\bra{\Omega_{\mu,\mu^{\prime},\mu^{\prime\prime}}}~{},$ (4.23)
where
$\bra{p_{x}^{\mu,\mu^{\prime},\mu^{\prime\prime}}}\in\bigotimes_{y~{}\mathrm{nn}~{}x}\mathcal{H}_{xy}\otimes\mathcal{H}_{\mu}^{*}\otimes\mathcal{H}_{\mu^{\prime}}^{*}\otimes\mathcal{H}^{*}_{\mu^{\prime\prime}}$ (4.24)
now includes some factors that are not associated to the gauge theory Hilbert space. A general state looks like
$\bra{p_{x}^{\mu,\mu^{\prime},\mu^{\prime\prime}}}=\sum_{k,k^{\prime},k^{\prime\prime}}P^{\mu\mu^{\prime}\mu^{\prime\prime}}_{kk^{\prime}k^{\prime\prime},\ell_{1}\ell_{2}\ell_{3}}\bra{\ell_{1}}_{xy_{1}}\cdots\bra{\ell_{3}}_{xy_{3}}\bra{\mu;k}\bra{\mu^{\prime};k^{\prime}}~{},\bra{\mu^{\prime\prime};k^{\prime\prime}}$ | (4.25)
and these $P^{\mu\mu^{\prime}\mu^{\prime\prime}}_{kk^{\prime}k^{\prime\prime},\ell_{1}\ell_{2}\ell_{3}}$ generally do not have an interpretation as a Clebsch-Gordan coefficient.
@@ 5 Conclusion
We have shown that there is a straightforward way to build holographic tensor networks that include geometries that are completely distinct, but nonetheless all must satisfy some constraints. As a byproduct, we found these tensor networks better relate to other ideas in holography, such as ensembles of random CFT data. This is a step towards adding time evolution in a way that resembles gravity. There, the constraints play a key role in having the bulk dynamics match those of the dual CFT. It will be interesting future work to incorporate different constraints, beyond just Gauss’ law, making the bulk theory more like gravity and the holographic map more realistic.

### v3

@@ Abstract
Conventional holographic tensor networks can be described as toy holographic maps constructed from many small linear maps acting in a spatially local way, all connected together with “background entanglement”, i.e. links of a fixed state, often the maximally entangled state. However, these constructions fall short of modeling real holographic maps. One reason is that their “areas” are trivial, taking the same value for all states, unlike in gravity where the geometry is dynamical. Recently, new constructions have ameliorated this issue by adding degrees of freedom that “live on the links”. This makes areas non-trivial, equal to the background entanglement piece plus a new positive piece that depends on the state of the link degrees of freedom. Nevertheless, this still has the downside that there is background entanglement, and hence it only models relatively limited code subspaces in which every area has a definite minimum value. In this note, we simply point out that a version of these constructions goes one step further: they can be background independent, with no background entanglement in the holographic map. This is advantageous because it allows tensor networks to model holographic maps for larger code subspaces. In addition to pointing this out, we address some subtleties involved in making it work.
@@ 1 Introduction
Let us summarize the network now. First, how might we make a tensor network without a rigid, fixed background? As a first guess, we might imagine a model in which we consider more than just one tensor network, allowing (somehow) for a “superposition of tensor networks”, each with a different geometry. This is a decent first step towards modeling gravity. However, it fails to model the fact that in gravity we do not allow arbitrary quantum states on a given geometry. There are constraints that the geometry must satisfy. These constraints are important and related to having good time evolution matching that of the dual theory. Really, we would like to construct a model akin to a “superposition of tensor networks” but in which the geometries of the tensor networks are required to satisfy some constraints.
@@
That said, these existing constructions arguably only go part of the way. They involve a background geometry, in a manner we’ll explain. The main goal of this note is to point out that a version of these constructions goes all the way, continuing to work even without the background geometry, once we have incorporated link degrees of freedom in the appropriate way. We explain this in Section 4. Thus we arrive at our goal: a tensor network incorporating multiple geometries, possibly all very distinct, and satisfying certain constraints. In Section 4.2 we address a subtlety in how to get this to work in one spatial dimension.
@@ 3 Link degrees of freedom
Now we turn to the holographic map. We would like a map composed of tensors acting in a spatially local way like in Section 2, but modified to act on this new $\mathcal{H}_{\mathrm{bulk}}$. The trick is we’ll first define a new kind of tensor that takes as input a link state in $\mathcal{H}_{G}$ and outputs a state in a factorized Hilbert space. Specifically, we define a tensor $S_{(xy)}:\mathcal{H}_{G}^{(xy)}\to\mathcal{H}_{G}^{xy}\otimes\mathcal{H}_{G}^{yx}$ to implement the following factorization (the usual embedding into the “extended Hilbert space” [23])
@@
with $\braket{\psi}{S^{\dagger}\hat{A}(b)S}{\psi}=\sum_{xy\in\partial b}\log D_{xy}$ (here for simplicity specializing to the case |$\ket{\phi_{xy}}=\ket{\mathrm{MAX}}$). This is an interesting answer, but we would like the answer for $V$ instead of $\widetilde{V}$. Fortunately, by construction $S(B)_{V\ket{\psi}}=S(B)_{\widetilde{V}S\ket{\psi}}$! All that’s different is the interpretation of the terms on the right hand side of (3.16). For $V$, the factors $\mathcal{H}_{G}^{xy}$ are not part of $\mathcal{H}_{\mathrm{bulk}}$, and so their entropy – the third term in the last line of (3.15) – is not a part of $S(b)_{\ket{\psi}}$. Instead, those terms must be interpretted as part of the area term. We end up with
@@ 4.2 $1d$ and background independence
Now let’s try to compute the entropy $S(B)$ for $B$ either of the two black circles. All bond dimensions are equal, so the minimal $b$ in the formula (4.10) can have many problems. For example, it can be maximally degenerate, with every cut an equally minimal surface, say if the qudits are in a product state. More problematically, if any bulk legs are in a mixed state, then neither boundary region will include that leg in its minimal $b$. (Though the entire boundary might still include the entire bulk in its minimal $b$.) This is not like what we expect in $1$-dimensional versions of AdS/CFT, where small amounts of bulk entropy are not enough to greatly change the position of the quantum extremal surface of the left or right boundary.
@@ 5 Conclusion
We have shown that there is a straightforward way to build holographic tensor networks that include geometries that are completely distinct, but nonetheless all must satisfy some constraints. This is a step towards adding time evolution in a way that resembles gravity. There, the constraints play a key role in having the bulk dynamics match those of the dual CFT. It will be interesting future work to incorporate different constraints, beyond just Gauss’ law, making the bulk theory more like gravity and the holographic map more realistic.